# Generating Building-Level Heat Demand Time Series by Combining Occupancy Simulations and Thermal Modeling

### v1

@@ Abstract
| INSPIRE | infrastructure for spatial information in the European community |
@@
| indoor temperature | |
| outside temperature | |
@@ 1 Introduction
While the general strategy described above seems conceptually simple, its real-world implementation presents significant challenges. Boosting the thermal renovation of buildings and retrofitting heating systems requires well-targeted policies [8]. Moreover, ongoing electrification of heating systems will increase the overall power demand, potentially leading to higher peak loads. This could necessitate grid reinforcements to ensure stability and additional expansion of renewable power production [9, 10].
@@
On the one hand, high spatial resolution of heating demand data, e.g. at the building level, allows for the identification of the most effective renovation measures and enables the design of well-targeted policies like subsidy programs. This high spatial resolution is also essential when assessing the economic viability of district heating networks [13]. On the other hand, high temporal resolution, e.g. hourly resolved heat demand time series, enables the analysis of the impact of heating system electrification on the power grid [14]. By considering the additional electrical loads (e.g. from heat pumps) and analyzing the impact of thermal load shifting on the power system, these data can inform decisions for sector-coupled generation and expansion planning (e.g. Ref. [15, 16]) in integrated power and heat systems. Together that shows that cost-efficient system planning and optimization strongly rely on granular data. This is often available (for research purposes) for the power sector as smart meter measurements at the building or household level, distribution network data, and national energy statistics, all with hourly or even quarter-hourly resolution. However, this level of detail is not yet available for heat load curves at any scale. Existing methodologies (discussed in detail in Section 1.1) and data sources for heating demand provide either geolocated annual heating demand, synthetic load profiles for individual buildings, or aggregated load curves for specific regions based on gas or district heating consumption.
@@ 1.1.1 Methods for spatial heating demand data
Beyond this standard approach, some models consider additional input parameters. For example, the Building Thermal Energy Assessment Model by Rades-Gil et al. [32] includes solar irradiation, which has been found to significantly impact specific energy demand, second only to climate variables [33]. Martín-Consuegra [29] also incorporates the facade area, allowing for higher spatial resolution even within building blocks.
@@ 1.1.2 Methods for temporal heat load profiles
Data-driven models require minimal information about the actual physical parameters of buildings. However, they do rely on measured heating demands, typically obtained from existing district heating networks or gas consumption data [45]. These models often aim to predict future consumption in existing district heating networks or perform correlation analyses between heating demand and external variables such as temperature. For example, Dang et al. [46, 47] analyzed heat load patterns in existing district heating systems, emphasizing the importance of considering not just thermal building parameters and outside temperature, but also user behavior for accurate predictions. Similarly, Maljkovic et al. [48] explored the influence of various parameters on heating demand using machine learning, measured data, and surveys. Fumo and Biswas [49] provide a review of regression analysis for predicting residential energy consumption, highlighting the simplicity of regression models compared to physical models for demand prediction. However, when no data is available, one must rely on engineering or physical models instead.
@@
At the database level, the When2Heat project [57, 58] provides demand data and other parameters, such as the coefficient of performance for heat pumps, for several European countries. This comprehensive database offers quick and easy access to valid load profiles. However, these data are aggregated at a x grid level, which does not provide sub-district level spatial resolution.
@@ 1.1.3 Combined approaches
Some studies offer spatially and temporally resolved data for specific regions, but not at the building level: Malla [19], in his master’s thesis, recognizes the need for high spatial and temporal resolution in heat demand profiles and presents a methodology to calculate load profiles for a x grid across Germany. Clegg [60] developed a method with a half-hourly resolution for 404 areas of Great Britain. While they validate their time series against measured gas consumption data, the spatial resolution remains at the district level. Lombardi et al. [12] present a thermodynamic modeling approach for Italian regions, incorporating many physical details for heat transfer to build an accurate model. However, the spatial resolution is limited to the NUTS-2 level. Their model (with available code) also enables running different refurbishment scenarios defined by the user.
@@
Finally, Heidenthaler [18] uses energy performance certificates and a simulation tool to generate load profiles at the building level. However, this model does not account for the stochastic nature of user behavior, leading to unrealistic peaks in heating demand.
@@ 1.2 Original contribution
In this paper, we present an integrated methodology that (i) leverages easily accessible, publicly available data, (ii) combines this data with a thermal model and stochastic active occupancy modeling, and (iii) generates hourly resolved time series at the building level for user-defined regions. This approach offers a straightforward solution for data generation, accounting for stochastic user behavior and the asynchronicity among different households. By providing the developed software, the method can be readily applied to generate data for a wide range of studies and investigations, thus addressing the methodological gaps highlighted in several studies [17, 59, 45].
@@ 2 Methodology
In this section, we present a methodology to generate hourly heat demand profiles on the single building level based on publicly accessible data and considering the stochasticity of user behavior. After discussing different data sources in 2.1, each step of the model is explained in detail in Subsections 2.2.1 - 2.2.6. Subsequently, Subsection 2.3 illustrates the validity of the method using several comparative and analytic approaches.
@@ 2.2.2 Step 2: Cadaster data request
The EU directive 2007/2/EG, passed in 2007 [62], aims to ensure the provision of comprehensive geospatial information across the EU. To achieve this, INSPIRE (Infrastructure for Spatial Information in the European Community) was established, enabling access to cadastral data in all member states. While the specific implementation varies by country and region, options such as bulk downloads or individual data requests based on address, cadastral reference, or GPS location are generally available. Ideally, these datasets provide information at the building level, including the number of units, usage type, year of construction, and (residential) area . If such data is unavailable, it can be approximated using alternative parameters, such as building shape area and height, as demonstrated in Ref. [25]. To apply our method, this information is retrieved from the cadastral records and appended to each building entry in the dataset.
@@ 2.2.4 Step 4: Estimation of thermal properties
Thermal losses, including transmission through walls, floors, roofs, and ventilation, are aggregated into a single parameter: the thermal conductance (in ). This is estimated by dividing the total annual heating demand by the cumulative indoor-outdoor temperature difference over all hours of the year.
@@ 2.2.5 Step 5: Synthetic load profile generation
The general distribution of active occupancy can be derived from Time Use Surveys [54], which are regularly conducted across European countries. These surveys capture daily routines through empirical questionnaires administered to a statistically representative sample population. However, applying a single, uniform time profile for all buildings in the thermal model would not accurately reflect real-world variability, as people do not wake up or follow their daily routines at the same time. To address this, the MCMC method [53] is used to generate individual active occupancy profiles for each dwelling . In an MCMC chain, the next state depends only on the current state and is determined by a transition matrix specific to the time step, with a randomness factor introduced for transitions. Different transition matrices are used for weekdays and weekends to reflect workday-dependent behavior. The randomness in the generated profiles captures individual variability in daily routines. Nevertheless, the average of a sufficiently large number of profiles converges to the distribution derived from TUS data.
Figure 3 illustrates this process, showing individual generated profiles (Subfigure 3(a)) and the average of many profiles (Subfigure 3(b)), converging for large sample numbers to the overall desired distribution.
@@
To calculate the actual current indoor temperature , the total heat provided by heating , plus internal gains , minus the thermal losses from the previous time step, is divided by the thermal capacity of the building and added to the previous indoor temperature (see Eq. (3)). Internal gains arise from using electrical appliances, cooking, etc., and are fixed to an average value. Initial values of and are chosen.
@@
Finally, the actual heating demand is determined by a case differentiation in Eq. (6): Negative heating demands (i.e., cooling demands) are set to zero. Additionally, a maximum heating power is introduced to limit the added heat per unit time, better representing a real system setup.
@@ 2.2.6 Step 6: Flexible data aggregation
The heating demand profiles obtained at the building level can now be directly applied for further data analysis, or incorporated into simulations or optimization models. Using one of the various methods for network generation, such as minimum spanning tree or other approaches [66, 25], a district heating network topology can be proposed. Based on this topology, the load profiles can be easily aggregated by summing all the profiles from a particular branch, thereby drastically reducing computational effort.
@@ 2.5 Methodical and data uncertainty
As with any modeling approach, the quality of input data is critical to the accuracy of the output. For heating demand at the building level, uncertainties arise from incomplete or outdated data in GIS/OSM or cadaster systems, as well as from ambiguities in the assignment of building archetypes, as discussed by Dochev et al. [67]. However, Dochev et al. also note that such discrepancies tend to average out when considering aggregated data. Potential systematic errors may also arise from neglecting additional factors, such as solar irradiation, wind chill, internal heat gains from the human body, lighting, cooking, other electrical appliances, or varying individual setpoint temperatures and ventilation habits. Nonetheless, Staffell et al. [33] demonstrate that, in their model, most of these factors have only a minor effect, with solar gain being the most significant. In comparison, the impact of these factors is generally smaller than the uncertainty introduced by the input data.
At this stage, we have intentionally opted for a simplified model, accepting some degree of uncertainty. It is worth noting that variations between different climate years may result in larger deviations than the factors discussed above.
@@ 3.2 Data analysis
For the Puertollano case study, the initial dataset for the region was sourced from OSM [23], which provides details on various infrastructures, including 11,725 buildings, of these, 9,298 are residential buildings. Each building object contains information about the centroid location, shape (defined by GPS coordinates), and height recorded.
Subsequent to this, additional building-specific data was obtained from the Spanish cadaster [70] using the provided API [70]. This dataset includes information on building usage type, construction year, area designation, number of units, and postal address. For buildings containing multiple units (e.g. apartment buildings), data for each individual subobject (household) was requested and aggregated at the building level. A sample of this data at the building level is presented in Figure 9.
@@
In the next step, the thermal load profile for each building was computed. The thermal properties were estimated using the available data, as outlined in Section 2.2.4. The MCMC method was then employed to generate an active occupancy profile, , with an hourly resolution for both business days and weekends. The transition matrices were derived from Ref. [53]. For the outdoor temperature data, the dataset from Rnewables.ninja [61] was utilized. These inputs were subsequently processed through the thermal model, where temperatures and heating demands were iteratively calculated over 8760 steps. This procedure resulted in an hourly heating demand time series for each building. The same process was also applied using data for the nominal heating demand of the buildings, assuming they were retrofitted. The generated dataset can be accessed in Ref. [71].

### v2

@@ Abstract
| INSPIRE | infrastructure for spatial information in the |
| European community | |
@@
| indoor temperature for time step | |
| outside temperature for time step | |
@@ 1 Introduction
While the general strategy described above seems conceptually simple, its real-world implementation presents significant challenges. Boosting the thermal renovation of buildings and retrofitting heating systems requires well-targeted policies [8]. Moreover, ongoing electrification of heating systems will increase the overall power demand, potentially leading to higher peak loads [9]. This could necessitate grid reinforcements to ensure stability and additional expansion of renewable power production [10].
@@
On the one hand, high spatial resolution of heating demand data, e.g. at the building level, allows for the identification of the most effective renovation measures and enables the design of well-targeted policies like subsidy programs. This high spatial resolution is also essential when assessing the economic viability of district heating networks [13]. On the other hand, high temporal resolution, e.g. hourly resolved heat demand time series, enables the analysis of the impact of heating system electrification on the power grid [14]. By considering the additional electrical loads (e.g. from heat pumps) and analyzing the impact of thermal load shifting on the power system, these data can inform decisions for sector-coupled generation and expansion planning in integrated power and heat systems. For example, focusing on the integration of district heating systems [15] or energy hubs [16]. Together that shows that cost-efficient system planning and optimization strongly rely on granular data. This is often available (for research purposes) for the power sector as smart meter measurements at the building or household level, distribution network data, and national energy statistics, all with hourly or even quarter-hourly resolution. However, this level of detail is not yet available for heat load curves at any scale. Existing methodologies (discussed in detail in Section 1.1) and data sources for heating demand provide either geolocated annual heating demand, synthetic load profiles for individual buildings, or aggregated load curves for specific regions based on gas or district heating consumption.
@@ 1.1.1 Methods for spatial heating demand data
Beyond this standard approach, some models consider additional input parameters. For example, the Building Thermal Energy Assessment Model by Prades-Gil et al. [32] includes solar irradiation, which has been found to significantly impact specific energy demand, second only to climate variables [33]. Martín-Consuegra [29] also incorporates the facade area, allowing for higher spatial resolution even within building blocks.
@@ 1.1.2 Methods for temporal heat load profiles
Data-driven models require minimal information about the actual physical parameters of buildings. However, they do rely on measured heating demands, typically obtained from existing district heating networks or gas consumption data [45]. These models often aim to predict future consumption in existing district heating networks or perform correlation analyses between heating demand and external variables such as temperature. For example, Dang et al. analyzed heat load patterns in existing district heating systems, emphasizing the importance of considering not just thermal building parameters and outside temperature [46], but also, in a subsequent work, user behavior for accurate predictions [47]. Similarly, Maljkovic and Basic [48] explored the influence of various parameters on heating demand using machine learning, measured data, and surveys. Fumo and Biswas [49] provide a review of regression analysis for predicting residential energy consumption, highlighting the simplicity of regression models compared to physical models for demand prediction. However, when no data is available, one must rely on engineering or physical models instead.
@@
At the database level, the When2Heat project (original publication [57], current version of the dataset [58]) provides demand data and other parameters, such as the coefficient of performance for heat pumps, for several European countries. This comprehensive database offers quick and easy access to valid load profiles. However, these data are aggregated at a x grid level, which does not provide sub-district level spatial resolution.
@@ 1.1.3 Combined approaches
Some studies offer spatially and temporally resolved data for specific regions, but not at the building level: Malla [19], in his master’s thesis, recognizes the need for high spatial and temporal resolution in heat demand profiles and presents a methodology to calculate load profiles for a x grid across Germany. Clegg and Mancarella [60] developed a method with a half-hourly resolution for 404 areas of Great Britain. While they validate their time series against measured gas consumption data, the spatial resolution remains at the district level. Lombardi et al. [12] present a thermodynamic modeling approach for Italian regions, incorporating many physical details for heat transfer to build an accurate model. However, the spatial resolution is limited to the NUTS-2 [61] level (corresponds to regions/provinces). Their model (with available code) also enables running different refurbishment scenarios defined by the user.
@@
Finally, Heidenthaler et al. [18] uses energy performance certificates and a simulation tool to generate load profiles at the building level. However, this model does not account for the stochastic nature of user behavior, leading to unrealistic peaks in heating demand.
@@ 1.2 Original contribution
The literature review highlights the variety of existing methods and approaches, each with a specific scope of application and inherent limitations. For instance, traditional building energy models can produce highly precise and, when properly calibrated, accurate heat demand time series for individual buildings or small building clusters. However, they require detailed input data – such as building geometry, orientation, wall structure, construction materials, and insulation characteristics – which may not be readily available for large-scale applications. In contrast, regression-based models can provide reasonable estimates at aggregated levels (e.g., district or city scale), but they typically lack the spatial granularity required to resolve heat demand at the individual building level.
To the best of our knowledge, none of the existing approaches can generate stochastic heat demand time series at the building level using only publicly available data for a large number of buildings. While individual features have been implemented into existing models, none of them combine all these elements into a unified framework. However, such integration is essential to enable the straightforward generation of heat demand data in data-scarce contexts, supporting a wide range of applications in energy system modeling – particularly those focusing on sector coupling with high spatial resolution, such as the planning of local heat networks for waste heat utilization. To address this methodological gap, we present a novel approach that integrates the following distinct features into a single workflow:
1.
Publicly Available Data Utilization – Tailoring the workflow to work exclusively with publicly available data ensures broad applicability and superior scalability, enabling analysis of thousands of buildings and overcoming data shortages.
2.
Simplified Physics-Based Thermal Model – Using a physics-based model allows for the generation of reasonable demand profiles at the individual building level, making the method independent of standard load profiles, disaggregation techniques, or previously measured profiles.
3.
Stochastic Occupancy Simulation – Incorporating user behavior and utilizing stochastic simulation ensures realistic demand profiles that account for household asynchronicity and demand fluctuations, while maintaining scalability from a single neighborhood to entire regions.
This novel approach enables the flexible generation of high-resolution heat demand data with minimal input requirements – supporting the decarbonization of heat supply across diverse regions. Although showcased through a specific case study, the method is widely applicable and highly scalable, offering a powerful tool for data-driven energy system planning.
@@ 2 Methodology
In this section, we present a methodology to generate hourly heat demand profiles on the single building level based on publicly accessible data and considering the stochasticity of user behavior. The process is designed to work with a minimum input of publicly available data, ensuring broad applicability. By that, it provides valuable data for various energy analyses across different fields of research, particularly where this data is currently lacking or unavailable.
After discussing different data sources in 2.1, each step of the model is explained in detail in Subsections 2.2.1 - 2.2.6. Subsequently, Subsection 2.3 illustrates the validity of the method using several comparative and analytic approaches.
@@ 2.2.2 Step 2: Cadaster data request
The EU directive 2007/2/EG, passed in 2007 [63], aims to ensure the provision of comprehensive geospatial information across the EU. To achieve this, INSPIRE (Infrastructure for Spatial Information in the European Community) was established, enabling access to cadastral data in all member states. While the specific implementation varies by country and region, options such as bulk downloads or individual data requests based on address, cadastral reference, or GPS location are generally available. Ideally, these datasets provide information at the building level, including the number of units, usage type, year of construction, and (residential) area . If such data is unavailable, it can be approximated using alternative parameters: The area can be estimated from the building’s shape and height, as demonstrated in Ref. [25]. For buildings with missing construction year data, a default yearly heating demand, representing a weighted average, is assigned in the subsequent steps. These fallback assumptions enhance the method’s robustness against missing data entries. However, each additional estimation introduces a potential source of uncertainty. To quantify this error margin, the Supplemental Material [65] compares results from cadastral data with the simplified estimation based solely on OSM, showing that the difference in yearly heating demand amounts to . To apply our method, this information is retrieved from the cadastral records and appended to each building entry in the dataset.
@@ 2.2.4 Step 4: Estimation of thermal properties
Thermal losses, including transmission through walls, floors, roofs, and ventilation, are represented by a single parameter for each building : the thermal conductance (in ). In the absence of more detailed data, this parameter is estimated by dividing the total annual heating demand for the respective building by the cumulative indoor-outdoor temperature difference over all hours of the year:
(1) -
@@ 2.2.5 Step 5: Synthetic load profile generation
The general distribution of active occupancy can be derived from Time Use Surveys [54], which are regularly conducted across European countries. These surveys capture daily routines through empirical questionnaires administered to a statistically representative sample population. However, applying a single, uniform time profile for all buildings in the thermal model would not accurately reflect real-world variability, as people do not wake up or follow their daily routines at the same time. To address this, the MCMC method [53] is used to generate individual active occupancy profiles for each dwelling . In an MCMC chain, the next state depends only on the current state and is determined by a transition matrix specific to the time step, with a randomness factor introduced for transitions. Different transition matrices are used for weekdays and weekends to reflect workday-dependent behavior. The randomness in the generated profiles captures individual variability in daily routines. Nevertheless, the average of a sufficiently large number of profiles converges to the distribution derived from TUS data. A detailed implementation of the MCMC method, including sample data and validation, can be found in the Supplementary Materials [65].
Figure 3 illustrates this process, showing individual generated profiles (Subfigure 3(a)) and the average of many profiles (Subfigure 3(b)), converging for large sample numbers to the overall desired distribution. It should be noted that due to the stochastic nature of MCMC, the individual profiles will differ with each simulation run. However, this does not impact the aggregated results. In cases where reproducibility is desired (e.g., to obtain consistent numbers for testing), each building can be assigned a seed number, to produce a reproducible series of random numbers.
@@
To calculate the actual current indoor temperature , the total heat provided by heating , plus internal gains and solar gains , minus the thermal losses from the previous time step, is divided by the thermal capacity of the building and added to the previous indoor temperature (see Eq. (4)). Internal gains arise from the use of electrical appliances, cooking, or the body heat of occupants. Depending on the available information and the required level of detail, a constant contribution can be assumed [45], the profile can be linked to active occupancy [52], or it can be derived from empirical measurements [68]. In our approach, the magnitude of internal gains per building is scaled by the number of dwellings. Solar gains account for the thermal energy received by the building from solar radiation. This contribution can be roughly estimated using hourly irradiation and window area, as discussed in [20] and Section 2.5. Initial values of and are chosen.
@@
Finally, the actual heating demand is determined by a case differentiation in Eq. (7). If the calculation results in a negative value for a particular hour, it is set to zero. Otherwise, a negative heating demand would imply a decrease in indoor temperature, which does not reflect real-world conditions. Additionally, a maximum heating power is introduced to limit the added heat per unit time, better representing a real system setup.
@@ 2.2.6 Step 6: Flexible data aggregation
The heating demand profiles obtained at the building level can now be directly applied for further data analysis, or incorporated into simulations or optimization models. Using one of the various methods for network generation, such as minimum spanning tree for previously clustered buildings [69] or at the individual building level [25], a district heating network topology can be proposed. Based on this topology, the load profiles can be easily aggregated by summing all the profiles from a particular branch, thereby drastically reducing computational effort.
@@
While the annual heating demand calculated by our method is deterministic for each individual building, the corresponding time series exhibit stochastic variations due to the MCMC-based occupancy modeling. By aggregating time series, individual patterns are smoothed out, leading to a convergence towards a common distribution. Since only the active occupancy profiles themselves have a stochastic component, it can be assumed that the stochastically influenced part of the heat profiles will converge in the same manner (i.e., by aggregating approximately 10 dwelling profiles, as shown in the Supplementary Material [65]). Any remaining differences in the aggregated heat demand time series are primarily caused by variations in the respective building stock.
@@ 2.5 Methodical and data uncertainty
As with any modeling approach, the quality of input data is critical to the accuracy of the output. For heating demand at the building level, uncertainties arise from incomplete or outdated data in GIS/OSM or cadaster systems, as well as from ambiguities in the assignment of building archetypes, as discussed by Dochev et al. [70]. However, Dochev et al. also note that such discrepancies tend to average out when considering aggregated data. For a specific application in a case study, we strongly encourage researchers applying this method to carefully examine the input data and conduct plausibility checks and sample validations.
Potential systematic errors may also arise from neglecting additional factors, such as solar irradiation, urban heat islands, wind chill, internal heat gains from the human body, lighting, cooking, other electrical appliances, or varying individual setpoint temperatures and ventilation habits. Nonetheless, Staffell et al. [33] demonstrate that, in their model, most of these factors have only a minor effect, with solar gain being the most significant. In comparison, the impact of these factors is generally smaller than the uncertainty introduced by the input data.
For projects that require precise solar gain calculations, more detailed building models are necessary. Accurately assessing shading effects and the amount of effective solar radiation reaching buildings requires consideration of building shapes, orientation, morphology, roof shapes and window openings [20]. Such analyses typically rely on highly detailed building data, including 3D models. These models are also crucial for evaluating the impact of urban heat islands [71], which, as reported by [72], can change the heating demand of individual buildings by up to 20%. However, comprehensive 3D building data is not yet widely accessible to the public. Therefore, we have deliberately chosen a simplified model, acknowledging a certain degree of uncertainty while ensuring broad and universal applicability of our approach. Beyond that, it is worth noting that variations between different climate years may result in larger deviations than the factors discussed above.
@@ 3 Case Study: Analyzing the heating demand of Puertollano/Spain
It is important to note that the applicability of the proposed method is not limited to this case study. This analysis illustrates the general structure of the obtained data and represents just one of many possible applications.
@@ 3.2 Data analysis
For the Puertollano case study, the initial dataset for the region was sourced from OSM [23], which provides details on various infrastructures, including buildings, of these, 9,298 are residential buildings. Each building object contains information about the centroid location, shape (defined by GPS coordinates), and height recorded.
Subsequent to this, additional building-specific data was obtained from the Spanish cadaster [75] using the provided API [76]. This dataset includes information on building usage type, construction year, area designation, number of units, and postal address. For buildings containing multiple units (e.g. apartment buildings), data for each individual subobject (household) was requested and aggregated at the building level. A sample of this data at the building level is presented in Figure 9. Due to the lack of detailed 3D building information and building-specific window areas, solar gains are not considered in this case study, such that in Equation (4). This simplification avoids introducing additional uncertainty from estimated or assumed solar exposure parameters, which would require complex geometric modeling and accurate material properties that are typically not available from open data sources.
@@
In the next step, the thermal load profile for each building was computed. The thermal properties were estimated using the available data, as outlined in Section 2.2.4. The MCMC method was then employed to generate an active occupancy profile, , with an hourly resolution for both business days and weekends. The transition matrices were derived from Ref. [53]. For the outdoor temperature data, the dataset from Renewables.ninja [62] was utilized. These inputs were subsequently processed through the thermal model, where temperatures and heating demands were iteratively calculated over 8760 steps. This procedure resulted in an hourly heating demand time series for each building. The same process was also applied using data for the nominal heating demand of the buildings, assuming they were retrofitted. The generated dataset can be accessed in Ref. [77]. The total computational time for calculating the time series for the city of buildings and 8760 hours is approximately on a standard laptop (Intel i7, 32 GB RAM). A detailed overview of the computational performance can be found in the Supplementary Material [65].